OpenStudy (anonymous):

Find the expression that represents the pattern and find 6th term of the pattern: 2, 7, 12, 17, 22...

myininaya (myininaya):

2+5=7 7+5=12 12+5=17 17+5=22 22+5=27 is the 6th term \[f_n=f_{n-1}+5\] \[ f_1=2\] so using this to find f_2 we do: \[f_2=f_{2-1}+5=f_1+5=2+5=7\] and so on...